# Ideals, Gröbner Bases, and PCPs

### v1

@@ Abstract
At the core of our new construction is a new class of alternatives to “sum-check” protocols. As used in past PCPs, these provide a method by which to verify that an $m$-variate degree $d$ polynomial $P$ evaluates to zero at every point of some set $S\subseteq\mathbb{F}_{q}^{m}$. Previous works had shown how to check this condition for sets of the form $S=H^{m}$ using $\mathcal{O}(m)$ queries with alphabet $\mathbb{F}_{q}^{d}$ assuming $d\geq|H|$. Our work improves this basic protocol in two ways: First we extend it to broader classes of sets $S$ (ones closer to Hamming balls rather than cubes). Second, it reduces the number of queries from $\mathcal{O}(m)$ to an absolute constant for the settings of $S$ we consider. Specifically when $S=(\{0,1\}^{m/c}_{\leq 1})^{c}$, where $T=\{0,1\}^{a}_{\leq b}\subseteq\mathbb{F}_{q}^{a}$ denotes the set of Boolean vectors of Hamming weight at most $b$ in $\mathbb{F}_{q}^{a}$, we give such an alternate to the sum-check protocol with $\mathcal{O}(1)$ queries with alphabet $\mathbb{F}_{q}^{\mathcal{O}(c+d)}$, using proofs of size $q^{\mathcal{O}(m^{2}/c)}$. Our new protocols use insights from the powerful theory of Gröbner bases to extend previously known protocols to these new settings with surprising ease. In doing so, they highlight why these theories from algebra may be of further use in complexity theory.
@@ 1 Introduction
In this paper, we give a new general framework for constructing algebraic PCPs that leads to the first proof of the PCP theorem using only one “composition” step. Starting with the work of Arora and Safra [AroraS], composition of PCPs has been a key ingredient in all previous PCP constructions. The original proof of the PCP theorem due to Arora, Lund, Motwani, Sudan, and Szegedy [ALMSS] used two composition steps, while the novel alternate proof due to Dinur [Dinur] uses $O(\log n)$ composition steps. Compositions improve various parameters of the construction at the cost of making the verifier less transparent. So it is a natural and long-sought goal to try to minimize the number of composition steps (to one, or even zero!). We achieve the weaker goal here.
@@
The key to our construction is a new class of protocols replacing the “sum-check” protocol in PCP constructions. The sum-check protocol, due to Lund, Fortnow, Karloff, and Nisan [LFKN] (also used in proofs of IP=PSPACE  [Shamir] and MIP=NEXPTIME [BFL]) has been a central ingredient in previous PCP constructions. In PCP constructions, the protocol is used to establish that an $m$-variate polynomial $P$ over $\mathbb{F}_{q}$ given as an oracle from $\mathbb{F}_{q}^{m}\to\mathbb{F}_{q}$ is identically zero on the set $\{0,1\}^{m}\subseteq\mathbb{F}_{q}^{m}$ or more generally, on some set of the form $H^{m}$ for $H\subseteq\mathbb{F}_{q}$. We refer to this latter task as “zero-on-variety testing”. (The reason for the use of the term “variety” to describe the set $H^{m}$ will become clearer later.) Other than the sum check protocol, the only other direct protocol for zero-on-variety testing is a protocol due to Ben-Sasson and Sudan [BenSassonS], which also only works for varieties of the form $H^{m}$. In this work, we establish a new connection between the theory of Gröbner bases and the zero-on-variety test of [BenSassonS] that allows us to get efficient zero-on-variety tests for a much broader class of varieties, including some varieties that are close to Hamming balls of constant radii. This latter setting which had eluded previous works and is key to our PCP construction.
@@
Armed with this new class of protocols, we show how to significantly simplify the ALMSS PCP construction. We start by giving a new PCP construction that works relative to any variety $V\subseteq\mathbb{F}_{q}^{m}$ with performance depending on the “Gröbner basis complexity” of the variety, a notion we define. We then show how to specialize this PCP in two different ways by using two different varieties — the first giving the usual $\mathcal{O}(\log n)$ randomness and $\operatorname{poly}(\log n)$ query PCP for $\mathsf{NP}$, and the second giving an $\mathcal{O}(n^{\varepsilon})$ randomness and $\mathcal{O}_{\varepsilon}(1)$ query PCP for NP, for any $\varepsilon>0$. No natural PCP (built without composition) was known with the latter setting of parameters, and indeed, this has been the key bottleneck in reducing the number of composition steps in ALMSS. We stress that both our ingredient PCPs are instantiations of the same protocol — only the choice of the variety is different (and the analysis of the Gröbner basis complexity of the varieties is straightforward)! And furthermore, our PCPs are already “robust assignment testers” in the sense of Dinur and Reingold [DinurReingold] (or equivalently, Robust PCPs of Proximity in the sense of Ben-Sasson, Goldreich, Harsha, Sudan and Vadhan [BGHSV]) and thus immediately composable. Putting our two robust assignment testers together yields our final PCP. The resulting proof thus gives the following simplifications to the ALMSS protocol: It eliminates one composition step, it eliminates the need for the “Hadamard PCP” entirely, and it eliminates the need for the “parallelization/robustification” step in ALMSS [ALMSS, Section 7].
@@ 1.1 Overview of Our Construction
Prior to this work, the only natural zero-on-variety tests considered the setting where $V=H^{m}$ for some $H\subseteq\mathbb{F}_{q}$. The protocol given by [BenSassonS] uses the following identity. $f\in\mathbb{I}(H^{m})$ if and only if there exist polynomials $f_{1},\ldots,f_{m}$ of degree at most d|H|d-|H| such that
@@
where $Z_{H}(Y)=\prod_{\alpha\in H}(Y-\alpha)$ is the canonical univariate polynomial that vanishes on $H$. The identity above follows from Alon’s Combinatorial Nullstellensatz [Alon:Null] and leads to a tester as follows: The zero-on-variety tester for $H^{m}$ expects oracles for $f_{1},\ldots,f_{m}$ as proof. It verifies using the low-degree test that each of these oracles has degree at most d|H|d-|H| and then verifies Equation 1 for a random choice of $X=(a_{1},\ldots,a_{m})$. Modulo further details (involving local correction), this leads to an $\mathcal{O}(m)$ local tester with randomness $\mathcal{O}(m\log q)$ and alphabet size $\mathcal{O}(d\log q)$.
@@
For our purposes, this choice of variety is insufficient. (Furthermore, the dependence of the locality on $m$ is also problematic, but we’ll address this later.) To remedy this, we use an alternate interpretation of the identity above. In this interpretation the identity holds because $\{Z_{H}(X_{1}),\ldots,Z_{H}(X_{m})\}$ form a “Gröbner basis” of the ideal $\mathbb{I}(H^{m})$ under a “graded monomial ordering”. We won’t define the exact notion of a Gröbner basis under different monomial orderings here — we don’t need to. The notion that suffices for us is a notion we call a Gröbner generating set $G$ of an ideal $\mathbb{I}$: $G$ is a Gröbner generating set for $\mathbb{I}$ if for all polynomials $P\in\mathbb{I}$ there exist polynomials $h_{g},g\in G$ such that $P=\sum_{g\in G}h_{g}\cdot g$ and $\deg(h_{g}\cdot g)\leq\deg(P)$ for every $g\in G$. And the above identity is the special case where $V=H^{m}$ with $\{Z_{H}(X_{1}),\ldots,Z_{H}(X_{m})\}$ as the Gröbner generating set. The zero-on-variety test of [BenSassonS] can now be extended to any variety that has a “small” Gröbner generating set. Indeed, this motivates our notion of the Gröbner complexity of a variety $V$ (see Definition 4.2 for a formal definition) — which is the size of the smallest Gröbner generating set $G$ of $\mathbb{I}(V)$. One crucial example for us is the following: The set of polynomials $\{X_{i}X_{j}|1\leq i<j\leq m\}\cup\{X_{i}^{2}-X_{i}|i\in[m]\}$ form a Gröbner generating set for the variety $\{0,1\}^{m}_{\leq 1}$ consisting of Boolean points of Hamming weight at most $1$ in $\mathbb{F}_{q}^{m}$. (Thus the variety $\{0,1\}^{m}_{\leq 1}$ has Gröbner complexity $\mathcal{O}(m^{2})$.) We also use the following basic property of Gröbner generating sets: If $G_{X}\subseteq\mathbb{F}_{q}[X]$ is a Gröbner generating set for variety $V_{1}$ and $G_{Y}\subseteq\mathbb{F}_{q}[Y]$ is one for $V_{2}$ then $G_{X}\cup G_{Y}\subseteq\mathbb{F}_{q}[X,Y]$ is a Gröbner generating set for $V_{1}\times V_{2}$. In particular this establishes that the Gröbner complexity of $(\{0,1\}^{m}_{\leq 1})^{c}\subseteq\mathbb{F}_{q}^{mc}$ is $\mathcal{O}(cm^{2})$. Applying the [BenSassonS] protocol to Gröbner generating sets now gives us a $\mathcal{O}(k)$ query protocol for testing zero-on-$V$ for variety $V$ of Gröbner complexity $k$.
@@
While this now gives many new varieties that have natural zero-on-variety tests, the locality of $\mathcal{O}(k)$ can be prohibitive. Our second contribution here is to give a new protocol to test this, that shifts the Gröbner complexity into the randomness of the protocol and achieves (a universal) constant locality. Specifically our verifier now expects an oracle for $\mathcal{M}_{V}(P)(X,Y)=\sum_{g\in G}Y_{g}h_{g}(X)$ where $Y=(Y_{g}|g\in G)$ is a new set of $k$ variables. Specifically performing a low-degree test on $\mathcal{M}_{V}(P)$ along with a test that verifies $M_{V}(P)(X,0)\equiv 0$ ensures that $\mathcal{M}_{V}(P)$ is effectively giving access to all linear combinations of $h_{g}(X)$ with constant query complexity. Testing the identity $P(X)=\sum_{g\in G}h_{g}(X)g(X)$ at a random choice of $X$ now requires only one query to $P$ and one to $M_{V}(P)$! Modulo some standard use of self-correction, this gives us an $\mathcal{O}(1)$ locality protocol for zero-on-$V$ testing with alphabet size $\mathcal{O}(d\log q)$ and randomness $\mathcal{O}((k+m)\log q)$ for any variety $V$ of Gröbner complexity $k$. (See Section 5.)
@@
For example, the 3-coloring protocol, based on a similar proof from [BenSassonS], goes as follows: Fix an odd prime power qq. [^2] We do this for simplicity here and allow us to assume $\{-1,0,1\}\subseteq\mathbb{F}_{q}$ can be used to represent $3$ distinct colors. The protocol easily extends to other fields using some other set of $3$ distinct elements of $\mathbb{F}_{q}$. For a variety $V\subseteq\mathbb{F}_{q}^{m}$ let its “extension degree” be the least integer $d$ such that every function $f:V\to\mathbb{F}_{q}$ can be extended to a degree $d$ polynomial in $\mathbb{F}_{q}[X_{1},\ldots,X_{m}]$. Now, given a variety $V\subseteq\mathbb{F}_{q}^{m}$ of Gröbner complexity $k$ and extension degree $d$ we consider the $3$-coloring problem on the vertex set $V$ (same $V$). Note that the graph is given by an edge function $E:V\times V\to\{0,1\}$ which can be shown to be extendable to a degree $2d$ polynomial $\hat{E}$ from $\mathbb{F}_{q}^{2m}\to\mathbb{F}_{q}$. A proof that $E$ is $3$-colorable includes polynomials $\chi:\mathbb{F}_{q}^{m}\to\mathbb{F}_{q}$, $A:\mathbb{F}_{q}^{m}\to\mathbb{F}_{q}$ and $B:\mathbb{F}_{q}^{2m}\to\mathbb{F}_{q}$ satisfying (1) $A(X)=\chi(X)\cdot(\chi(X)-1)\cdot(\chi(X)+1)$, (2) $A$ is zero-on-$V$, (3) $B(X,Y)=\hat{E}(X,Y)\prod_{i\in\{-2,-1,1,2\}}(\chi(X)-\chi(Y)-i)$ and (4) $B$ is zero on $V\times V$. (Items (1) and (2) verify that $\chi$ is a $3$-coloring of $V$ with color set $\{-1,0,1\}$, while items (3) and (4) verify that $\chi$ is a valid coloring of the edges of $E$.) The $V$-verifier performs low-degree tests on all the four oracles and then tests identities (1) and (3) by picking a random value of the variables, and finally verifies items (2) and (4) using zero-on-$V$ and zero-on-$V^{2}$ tests. By the aforementioned properties on Gröbner complexity and standard facts about extension degree we get that this PCP verifier achieves $\mathcal{O}(1)$ locality with randomness $\mathcal{O}((k+m)\log q)$ and alphabet $\mathcal{O}(d\log q)$ (matching those of the zero-on-$V$ tests up to constant factors).
@@
Instantiating the verifier above with $V=H^{m}$ where $|V|=n$, $|H|=\log n$, $m=\mathcal{O}(\log n/\log\log n)$ and $q=\mathcal{O}(\log^{6}n)$ gives an $\mathcal{O}(1)$ locality PCP verifier for 3-coloring of $n$ vertex graphs with randomness $\mathcal{O}(\log n)$ and alphabet size $\mathcal{O}(\operatorname{poly}\log n)$. But a different instantiation of the same PCP with $m=n^{\varepsilon}$, $c=\frac{1}{\varepsilon}$, $q=\mathcal{O}_{\varepsilon}(1)$ and $V=(\{0,1\}^{m}_{\leq 1})^{c}$ gives an $\mathcal{O}(1)$ locality PCP verifier with randomness $O(n^{2\epsilon})$ and alphabet size $O_{\epsilon}(1)$! We note that even using $c=1$ gives a completely new protocol matching the parameters of the Hadamard PCP in [ALMSS, Section 5]. And using larger values of $c$ gives us our new protocols. (See Section 6 for details.)
@@
Furthermore, these PCPs are easily converted to “Robust PCPs of Proximity” (or “Robust assignment testers”) in the sense of [DinurReingold, BGHSV] of constant robustness — since our PCPs have constant locality. This allows us to compose the PCPs above in a single composition step to get an $\mathcal{O}(1)$ locality PCP verifier with $\mathcal{O}(\log n)$ randomness and $\mathcal{O}(1)$ alphabet size — and thus the PCP theorem. (See Section 7 for details.)

### v2

# Ideals, Macaulay Bases, and PCPs

@@ Abstract
At the core of our new construction is a new class of alternatives to “sum-check” protocols. As used in past PCPs, these provide a method by which to verify that an $m$-variate degree $d$ polynomial $P$ evaluates to zero at every point of some set $S\subseteq\mathbb{F}_{q}^{m}$. Previous works had shown how to check this condition for sets of the form $S=H^{m}$ using $\mathcal{O}(m)$ queries with alphabet $\mathbb{F}_{q}^{d}$ assuming $d\geq|H|$. Our work improves this basic protocol in two ways: First we extend it to broader classes of sets $S$ (ones closer to Hamming balls rather than cubes). Second, it reduces the number of queries from $\mathcal{O}(m)$ to an absolute constant for the settings of $S$ we consider. Specifically when $S=(\{0,1\}^{m/c}_{\leq 1})^{c}$, where $T=\{0,1\}^{a}_{\leq b}\subseteq\mathbb{F}_{q}^{a}$ denotes the set of Boolean vectors of Hamming weight at most $b$ in $\mathbb{F}_{q}^{a}$, we give such an alternate to the sum-check protocol with $\mathcal{O}(1)$ queries with alphabet $\mathbb{F}_{q}^{\mathcal{O}(c+d)}$, using proofs of size $q^{\mathcal{O}(m^{2}/c)}$. Our new protocols use insights from the powerful theory of Macaulay bases to extend previously known protocols to these new settings with surprising ease. In doing so, they highlight why these theories from algebra may be of further use in complexity theory.
@@ 1 Introduction
In this paper, we give a new general framework for constructing algebraic PCPs that leads to the first proof of the PCP theorem using only one “composition” step. Starting with the work of Arora and Safra [3], composition of PCPs has been a key ingredient in all previous PCP constructions. The original proof of the PCP theorem due to Arora, Lund, Motwani, Sudan, and Szegedy [2] used two composition steps, while the novel alternate proof due to Dinur [10] uses $O(\log n)$ composition steps. Compositions improve various parameters of the construction at the cost of making the verifier less transparent. So it is a natural and long-sought goal to try to minimize the number of composition steps (to one, or even zero!). We achieve the weaker goal here.
@@
The key to our construction is a new class of protocols replacing the “sum-check” protocol in PCP constructions. The sum-check protocol, due to Lund, Fortnow, Karloff, and Nisan [15] (also used in proofs of IP=PSPACE  [21] and MIP=NEXPTIME [4]) has been a central ingredient in previous PCP constructions. In PCP constructions, the protocol is used to establish that an $m$-variate polynomial $P$ over $\mathbb{F}_{q}$ given as an oracle from $\mathbb{F}_{q}^{m}\to\mathbb{F}_{q}$ is identically zero on the set $\{0,1\}^{m}\subseteq\mathbb{F}_{q}^{m}$ or more generally, on some set of the form $H^{m}$ for $H\subseteq\mathbb{F}_{q}$. We refer to this latter task as “zero-on-variety testing”. (The reason for the use of the term “variety” to describe the set $H^{m}$ will become clearer later.) Other than the sum check protocol, the only other direct protocol for zero-on-variety testing is a protocol due to Ben-Sasson and Sudan [6], which also only works for varieties of the form $H^{m}$. In this work, we establish a new connection between the theory of Macaulay bases and the zero-on-variety test of [6] that allows us to get efficient zero-on-variety tests for a much broader class of varieties, including some varieties that are close to Hamming balls of constant radii. This latter setting which had eluded previous works and is key to our PCP construction.
@@
Armed with this new class of protocols, we show how to significantly simplify the ALMSS PCP construction. We start by giving a new PCP construction that works relative to any variety $V\subseteq\mathbb{F}_{q}^{m}$ with performance depending on a notion of Macaulay complexity of the variety that we define. It is based on the notion of H-bases (or Macaulay bases, and is related to the notion of Gröbner bases) which go all the way back to the work of Macaulay in the 1900s, see e.g. [19, 16, 7, 17]. We then show how to specialize this PCP in two different ways by using two different varieties — the first giving the usual $\mathcal{O}(\log n)$ randomness and $\operatorname{poly}(\log n)$ query PCP for $\mathsf{NP}$, and the second giving an $\mathcal{O}(n^{\varepsilon})$ randomness and $\mathcal{O}_{\varepsilon}(1)$ query PCP for NP, for any $\varepsilon>0$. No natural PCP (built without composition) was known with the latter setting of parameters, and indeed, this has been the key bottleneck in reducing the number of composition steps in ALMSS. We stress that both our ingredient PCPs are instantiations of the same protocol — only the choice of the variety is different (and the analysis of the Macaulay complexity of the varieties is straightforward)! And furthermore, our PCPs are already “robust assignment testers” in the sense of Dinur and Reingold [9] (or equivalently, Robust PCPs of Proximity in the sense of Ben-Sasson, Goldreich, Harsha, Sudan and Vadhan [5]) and thus immediately composable. Putting our two robust assignment testers together yields our final PCP. The resulting proof thus gives the following simplifications to the ALMSS protocol: It eliminates one composition step, it eliminates the need for the “Hadamard PCP” entirely, and it eliminates the need for the “parallelization/robustification” step in ALMSS [2, Section 7].
@@ 1.1 Overview of Our Construction
Prior to this work, the only natural zero-on-variety tests considered the setting where $V=H^{m}$ for some $H\subseteq\mathbb{F}_{q}$. The protocol given by [6] uses the following identity. $f\in\mathbb{I}(H^{m})$ if and only if there exist polynomials $f_{1},\ldots,f_{m}$ of degree at most d|H|d-|H| such that
@@
where $Z_{H}(Y)=\prod_{\alpha\in H}(Y-\alpha)$ is the canonical univariate polynomial that vanishes on $H$. The identity above follows from Alon’s Combinatorial Nullstellensatz [1] and leads to a tester as follows: The zero-on-variety tester for $H^{m}$ expects oracles for $f_{1},\ldots,f_{m}$ as proof. It verifies using the low-degree test that each of these oracles has degree at most d|H|d-|H| and then verifies Equation 1 for a random choice of $X=(a_{1},\ldots,a_{m})$. Modulo further details (involving local correction), this leads to an $\mathcal{O}(m)$ local tester with randomness $\mathcal{O}(m\log q)$ and alphabet size $\mathcal{O}(d\log q)$.
@@
For our purposes, this choice of variety is insufficient. (Furthermore, the dependence of the locality on $m$ is also problematic, but we’ll address this later.) To remedy this, we use an alternate interpretation of the identity above. In this interpretation the identity holds because $\{Z_{H}(X_{1}),\ldots,Z_{H}(X_{m})\}$ form a “Gröbner basis” of the ideal $\mathbb{I}(H^{m})$ under a “graded monomial ordering”. We won’t define the exact notion of a Gröbner basis under different monomial orderings here — we don’t need to. The notion that suffices for us is the notion of a Macaulay basis $\mathfrak{G}$ of an ideal $\mathbb{I}$: $\mathfrak{G}$ is a Macaulay basis for $\mathbb{I}$ if for all polynomials $P\in\mathbb{I}$ there exist polynomials $h_{g},g\in\mathfrak{G}$ such that $P=\sum_{g\in\mathfrak{G}}h_{g}\cdot g$ and $\deg(h_{g}\cdot g)\leq\deg(P)$ for every $g\in\mathfrak{G}$. And the above identity is the special case where $V=H^{m}$ with $\{Z_{H}(X_{1}),\ldots,Z_{H}(X_{m})\}$ as the Macaulay basis. The zero-on-variety test of [6] can now be extended to any variety that has a “small” Macaulay basis. Indeed, this motivates our notion of the Macaulay complexity of a variety $V$ (see Definition 4.2 for a formal definition) — which is the size of the smallest Macaulay basis $\mathfrak{G}$ of $\mathbb{I}(V)$. One crucial example for us is the following: The set of polynomials $\{X_{i}X_{j}|1\leq i<j\leq m\}\cup\{X_{i}^{2}-X_{i}|i\in[m]\}$ form a Macaulay basis for the variety $\{0,1\}^{m}_{\leq 1}$ consisting of Boolean points of Hamming weight at most $1$ in $\mathbb{F}_{q}^{m}$. (Thus the variety $\{0,1\}^{m}_{\leq 1}$ has Macaulay complexity $\mathcal{O}(m^{2})$.) We also use the following basic property of Macaulay bases: If $\mathfrak{G}_{X}\subseteq\mathbb{F}_{q}[X]$ is a Macaulay basis for variety $V_{1}$ and $\mathfrak{G}_{Y}\subseteq\mathbb{F}_{q}[Y]$ is one for $V_{2}$ then $\mathfrak{G}_{X}\cup\mathfrak{G}_{Y}\subseteq\mathbb{F}_{q}[X,Y]$ is a Macaulay basis for $V_{1}\times V_{2}$. In particular this establishes that the Macaulay complexity of $(\{0,1\}^{m}_{\leq 1})^{c}\subseteq\mathbb{F}_{q}^{mc}$ is $\mathcal{O}(cm^{2})$. Applying the [6] protocol to Macaulay bases now gives us a $\mathcal{O}(k)$ query protocol for testing zero-on-$V$ for variety $V$ of Macaulay complexity $k$.
@@
While this now gives many new varieties that have natural zero-on-variety tests, the locality of $\mathcal{O}(k)$ can be prohibitive. Our second contribution here is to give a new protocol to test this, that shifts the Macaulay complexity into the randomness of the protocol and achieves (a universal) constant locality. Specifically our verifier now expects an oracle for $\mathcal{M}_{V}(P)(X,Y)=\sum_{g\in\mathfrak{G}}Y_{g}h_{g}(X)$ where $Y=(Y_{g}|g\in\mathfrak{G})$ is a new set of $k$ variables. Specifically performing a low-degree test on $\mathcal{M}_{V}(P)$ along with a test that verifies $\mathcal{M}_{V}(P)(X,0)\equiv 0$ ensures that $\mathcal{M}_{V}(P)$ is effectively giving access to all linear combinations of $h_{g}(X)$ with constant query complexity. Testing the identity $P(X)=\sum_{g\in\mathfrak{G}}h_{g}(X)g(X)$ at a random choice of $X$ now requires only one query to $P$ and one to $\mathcal{M}_{V}(P)$! Modulo some standard use of self-correction, this gives us an $\mathcal{O}(1)$ locality protocol for zero-on-$V$ testing with alphabet size $\mathcal{O}(d\log q)$ and randomness $\mathcal{O}((k+m)\log q)$ for any variety $V$ of Macaulay complexity $k$. (See Section 5.)
@@
For example, the 3-coloring protocol, based on a similar proof from [6], goes as follows: Fix an odd prime power qq. [^2] We do this for simplicity here and allow us to assume $\{-1,0,1\}\subseteq\mathbb{F}_{q}$ can be used to represent $3$ distinct colors. The protocol easily extends to other fields using some other set of $3$ distinct elements of $\mathbb{F}_{q}$. For a variety $V\subseteq\mathbb{F}_{q}^{m}$ let its “extension degree” be the least integer $d$ such that every function $f:V\to\mathbb{F}_{q}$ can be extended to a degree $d$ polynomial in $\mathbb{F}_{q}[X_{1},\ldots,X_{m}]$. Now, given a variety $V\subseteq\mathbb{F}_{q}^{m}$ of Macaulay complexity $k$ and extension degree $d$ we consider the $3$-coloring problem on the vertex set $V$ (same $V$). Note that the graph is given by an edge function $E:V\times V\to\{0,1\}$ which can be shown to be extendable to a degree $2d$ polynomial $\hat{E}$ from $\mathbb{F}_{q}^{2m}\to\mathbb{F}_{q}$. A proof that $E$ is $3$-colorable includes polynomials $\chi:\mathbb{F}_{q}^{m}\to\mathbb{F}_{q}$, $A:\mathbb{F}_{q}^{m}\to\mathbb{F}_{q}$ and $B:\mathbb{F}_{q}^{2m}\to\mathbb{F}_{q}$ satisfying (1) $A(X)=\chi(X)\cdot(\chi(X)-1)\cdot(\chi(X)+1)$, (2) $A$ is zero-on-$V$, (3) $B(X,Y)=\hat{E}(X,Y)\prod_{i\in\{-2,-1,1,2\}}(\chi(X)-\chi(Y)-i)$ and (4) $B$ is zero on $V\times V$. (Items (1) and (2) verify that $\chi$ is a $3$-coloring of $V$ with color set $\{-1,0,1\}$, while items (3) and (4) verify that $\chi$ is a valid coloring of the edges of $E$.) The $V$-verifier performs low-degree tests on all the four oracles and then tests identities (1) and (3) by picking a random value of the variables, and finally verifies items (2) and (4) using zero-on-$V$ and zero-on-$V^{2}$ tests. By the aforementioned properties of Macaulay complexity and standard facts about extension degree we get that this PCP verifier achieves $\mathcal{O}(1)$ locality with randomness $\mathcal{O}((k+m)\log q)$ and alphabet $\mathcal{O}(d\log q)$ (matching those of the zero-on-$V$ tests up to constant factors).
@@
Instantiating the verifier above with $V=H^{m}$ where $|V|=n$, $|H|=\log n$, $m=\mathcal{O}(\log n/\log\log n)$ and $q=\mathcal{O}(\log^{6}n)$ gives an $\mathcal{O}(1)$ locality PCP verifier for 3-coloring of $n$ vertex graphs with randomness $\mathcal{O}(\log n)$ and alphabet size $\mathcal{O}(\operatorname{poly}\log n)$. But a different instantiation of the same PCP with $m=n^{\varepsilon}$, $c=\frac{1}{\varepsilon}$, $q=\mathcal{O}_{\varepsilon}(1)$ and $V=(\{0,1\}^{m}_{\leq 1})^{c}$ gives an $\mathcal{O}(1)$ locality PCP verifier with randomness $\mathcal{O}(n^{2\epsilon})$ and alphabet size $O_{\epsilon}(1)$! We note that even using $c=1$ gives a completely new protocol matching the parameters of the Hadamard PCP in [2, Section 5]. And using larger values of $c$ gives us our new protocols. (See Section 6 for details.)
@@
Furthermore, these PCPs are easily converted to “Robust PCPs of Proximity” (or “Robust assignment testers”) in the sense of [9, 5] of constant robustness — since our PCPs have constant locality. This allows us to compose the PCPs above in a single composition step to get an $\mathcal{O}(1)$ locality PCP verifier with $\mathcal{O}(\log n)$ randomness and $\mathcal{O}(1)$ alphabet size — and thus the PCP theorem. (See Section 7 for details.)
Related constructions.
Here we will discuss two PCP constructions that are relevant to us; both of them attain the parameters of our inner verifiers (i.e., $\mathcal{O}(1)$ locality and $\mathcal{O}(n^{\varepsilon})$ randomness, and $\mathcal{O}_{\varepsilon}(1)$ alphabet size) using constructions that don’t make use of a “full composition” but do perform certain transformations/encodings of an initial (sub-optimal) proof to achieve the final parameters.
First, we mention the work of Goldreich [12] posted shortly after our paper, which gives a different construction of a PCP matching our inner verifier parameters. This construction is based on using a Reed-Muller PCP at an outer level and encoding the larger field elements using a Hadamard code. The tests of the verifier then resemble that of the usual Hadamard PCP of [2]. Composing this inner verifier (rather the PCP of Proximity version of it) with the the (robust version of the) outer verifier of [2] also results in the PCP Theorem (in 1.5 proof compositions as Goldreich refers to it).
Second, we mention the work of [14] studying quasi-random PCPs, which have a very different behavior from the usual PCPs. While [14] obtains a constant query quasi-random PCP of size similar to our inner verifier (i.e., $2^{n^{\varepsilon}}$), the construction (among many other differences) involves using the Hadamard encodings of the elements of a large field. We also note that both the works of [14] as well as [12] also deviate from our techniques in that we do not perform a sum-check protocol in order to do the zero-on-variety tests.